### v2

# Landscape of Correlated Orders in Strained Bilayer Nickelate Thin Films

Abstract
The discovery of high-temperature superconductivity in bilayer nickelates La3Ni2O7 under pressure has sparked significant research interest. This interest has been further fueled by the recent achievement of superconductivity in compressed thin films at ambient pressure, although the origin and underlying mechanism remain elusive. In this work, we explore the electronic structures and instabilities of strained thin films on substrates to identify the key factors for achieving superconductivity, using first-principles and functional renormalization group (FRG) calculations. Our findings suggest that the compressed NiO2 bilayer at the interface is unlikely to exhibit superconductivity due to the elongation of outer apical Ni-O bond length and electron-doping effect. In contrast, the NiO2 bilayer slightly away from the interface shows density-wave instability when undoped or slightly hole-doped. However, when this bilayer is moderately hole-doped, leading to the emergence of a hole pocket around the M point, density-wave instability is suppressed and the system exhibits robust $s_{\pm}$-wave superconductivity, which may account for superconductivity observed in thin films. For the stretched NiO2 bilayer, robust spin-density-wave instability is observed due to enhanced Fermi surface nesting, despite the presence of a hole pocket around the M point. Potential experimental implications are discussed. Our study highlights the crucial role of fermiology in determining electronic instability and establishes a unified scenario for superconductivity in both pressurized bulk and strained thin films of bilayer nickelates.
Despite decades of extensive research efforts into nickelates, the nickel age of superconductors only began in 2019 with the discovery of superconductivity in ”infinite-layer” nickelates, specifically (Sr,Nd)NiO2 thin films grown on a substrate Li et al. (2019); Osada et al. (2020); Pan et al. (2022); Wang et al. (2022); Ding et al. (2023). Remarkably, four years later, nickelate superconductors entered the high-temperature regime akin to cuprates with the observation of superconductivity in a novel type of bilayer nickelate La3Ni2O7 (LNO) within the Ruddlesden-Popper phase under pressure, boasting an extraordinarily high transition temperature (Tc) of approximately 80 K Sun et al. (2023). Superconductivity emerges abruptly following a structural transition from the Amam phase to the Fmmm/I4mmm phase as pressure increases. In contrast to the $d^{9}$ electronic configuration of cuprates and ”infinite-layer” nickelates, the Ni2.5+ in LNO exhibits a $d^{7.5}$ configuration, with both $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals contributing to the low-energy electronic structure due to the presence of apical oxygens between layers Luo et al. (2023); Zhang et al. (2023a); Lechermann et al. (2023a); Sakakibara et al. (2024a); Gu et al. (2025); Yang et al. (2024); Liu et al. (2024). Additionally, the trilayer nickelate La4Ni3O10 exhibits superconductivity under pressure with a lower Tc of 20-30 K Sakakibara et al. (2024b); Li et al. (2024a); Zhu et al. (2024); Zhang et al. (2025); Li et al. (2024b). Both bilayer and trilayer nickelates have displayed multiple correlated phenomena, such as density wave transitions and non-Fermi liquid behaviors Sun et al. (2023); Zhang et al. (2024a); Liu et al. (2024); Chen et al. (2024); Ren et al. (2025). The pairing mechanism and symmetry are under intensive theoretical study Yang et al. (2023a); Lu et al. (2024); Oh and Zhang (2023); Gu et al. (2025); Liu et al. (2023); Qu et al. (2024); Sakakibara et al. (2024a); Yang et al. (2023b); Qin and Yang (2023); Lu et al. (2023); Tian et al. (2024); Zhang et al. (2023b, 2024b); Jiang et al. (2024a); Lechermann et al. (2023b); Liao et al. (2023); Ryee et al. (2024); Luo et al. (2024); Fan et al. (2024); Jiang et al. (2024b, c); Zhan et al. (2025); Xia et al. (2025); Bejas et al. (2025) and calculations suggest that the pressure-driven Lifshitz transition, where the $d_{z^{2}}$ interlayer bonding state crosses the Fermi level to form a hole pocket, is crucial for the emergence of bulk superconductivity Sun et al. (2023). Spin fluctuations are believed to promote an $s_{\pm}$-wave pairing following the Lifshitz transition Sakakibara et al. (2024a); Gu et al. (2025); Yang et al. (2023a); Liu et al. (2023). Other studies emphasize the importance of interlayer exchange coupling, Hund’s rule coupling, and orbital hybridization in generating interlayer $s$-wave pairing Lu et al. (2024); Oh and Zhang (2023); Qu et al. (2024). However, further experimental investigation into the electronic structures and pairing gaps in the bulk nickelates is challenged due to the high-pressure conditions required for superconductivity to occur.
Intriguingly, recent studies have shown that compressively strained LNO thin films grown on the SrLaAlO4 (SLAO) substrate exhibit superconductivity with a Tc exceeding 40 K at ambient pressure Ko et al. (2025); Zhou et al. (2025); Liu et al. (2025); Bhatt et al. (2025). Although this Tc is only about half that of the bulk value, it provides a tunable and powerful platform for investigating the mechanisms underlying high-Tc superconductivity. First-principles calculations indicate that compressive strain eliminates the hole pocket around the M point, whereas tensile strain results in Fermi surfaces similar to those of the bulk material under pressure Geisler et al. (2025); Zhao and Botana (2025). However, superconductivity has not been observed in tensile-strained LNO thin films on substrates Ko et al. (2025); Bhatt et al. (2025), such as SrTiO3, which seems to contradict the proposed mechanism for superconductivity in the bulk. Recent ARPES measurements have revealed heavy hole doping in La2.85Pr0.15Ni2O7 ultrathin films, identifying a diffusive hole pocket around the M point Li et al. (2025); Yue et al. (2025). These findings motivate a comprehensive exploration of the key ingredients for superconductivity in bilayer nickelates and prompt several key questions regarding LNO thin films: (1) Which bilayer is responsible for superconductivity in LNO thin films, and what is the underlying mechanism? (2) Why do tensile-strained LNO thin films, despite hosting Fermi surfaces similar to the bulk, fail to exhibit superconductivity? (3) Do bulk and thin film LNO share the same pairing mechanism?
To address these questions, we examine the electronic structures of LNO thin films on SLAO substrate and explore competing correlated instabilities, such as superconductivity and density-wave states, for NiO2 bilayers at different locations and under different in-plane strains. We employ first-principles and functional renormalization group (FRG) calculations to conduct our analysis. Our findings reveal that the NiO2 bilayer at the interface is heavily electron-doped due to the SLAO substrate and exhibits asymmetry between two Ni sites due to the elongation of outer apical Ni-O bond at the interface. This causes a significant downshift in the onsite energy of the $d_{z^{2}}$ orbital. Our FRG analysis indicates that the instability in this NiO2 bilayer manifests as a spin density wave (SDW) over a broad range of interaction parameters due to Fermi surface nesting, suggesting that superconductivity at the interface is unlikely. For the NiO2 bilayer slightly away from the interface, an SDW is still favored when it is undoped or slightly hole-doped. However, when this bilayer is moderately hole-doped, leading to a Lifshitz transition around the M point and the appearance of a hole pocket, additional Fermi surface nesting occurs, and it results in strong competition between two SDW orders, destabilizing the SDW instability. Remarkably, both Fermi surfaces nesting occur between bonding and antibonding pockets, resulting in interlayer antiferromagnetic coupling in both SDWs. Thus, their fluctuations promote interlayer pairing, leading to a robust $s_{\pm}$-wave superconductivity. This may account for the superconductivity observed in compressed thin films. Additionally, we find that tensile strain significantly enhances Fermi surface nesting between the $\alpha$ and $\beta$ pockets as well as the $\alpha$ and $\gamma$ pockets, inducing an SDW instability despite the presence of the hole pocket around the M point. We discuss a unified scenario for superconductivity in both bulk and thin films and explore potential experimental implications.
[…]
Electronic structures of LNO/SLAO interface. We start with the crystal and electronic structures of three-unit-cell (3UC) LNO on the SLAO substrate. The tetragonal LNO is assumed and initial interface is constructed according to the structure from scanning transmission electron microscope measurements Zhou et al. (2025); Yue et al. (2025) and further relaxation is done for the 3UC LNO and the first two layers of substrate. The details of first-principles calculations are shown in the supplementary materials (SM). Fig. 1(a) shows the optimized crystal structure, where only three bilayers of the film are shown along with part of the substrate, and the complete structural diagram is available in SM. The in-plane compressive strain from the substrate will generate an elongation along the c-axis for LNO, which mainly occurs on the LaO layers according to our calculations. To clearly demonstrate structural changes, we select the first and third bilayers as representative examples and show their structural parameters. The out-of-plane distance between two Ni sites $d_{\text{Ni}-\text{Ni}}$ and outer apical Ni-O bond length $d_{\text{Ni}-\text{O}}$ within these bilayers are indicated in gray and red numbers, respectively. Across these bilayers, $d_{\text{Ni}-\text{Ni}}$ is approximately 4 Å, close to the bulk value, and this value is slightly longer in the 1st bilayer than the 3rd bilayer due to the effect of substrate. Notably, the outer apical Ni-O distance is increased to 2.3 Å  away from the interface due to the elongation of LaO layers and $d_{\text{Ni}-\text{O}}$ in the 1st bilayer is elongated to 2.6 Å. This elongation of the outer apical Ni-O bond at the interface results in a significant reduction in the onsite energy of the $d_{z^{2}}$ orbital in the 1st bilayer. Besides structural modifications, there is electrostatic doping from the substrate. Fig. 1(b) displays charge density differences and blue and red indicate regions of charge accumulation and depletion, respectively, suggesting a net electron transfer from the substrate to the first bilayer.
Furthermore, the orbital-resolved band structures along high-symmetry paths for the 1st and 3rd bilayers are presented in Fig. 1(c) and (d), with a focus on the $e_{g}$ orbitals ($d_{x^{2}-y^{2}}$ and $d_{z^{2}}$), as the fully occupied $t_{2g}$ orbitals lie well below the Fermi level. In the 1st bilayer, the reduced onsite energy of the $d_{z^{2}}$ orbital induces a significant downward shift of its associated bands, as seen in Fig. 1(d). This shift causes the $d_{z^{2}}$ bonding band to move further from the Fermi level, while the $d_{z^{2}}$ anti-bonding states approach and make contact with the Fermi level at the $\Gamma$ point. The electron doping of this bilayer is estimated to be about 0.34. As a result, the 1st bilayer hosts only one $\beta$ hole pocket around the M point, contributed by the mixture of $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals, and an $\alpha$ electron pocket around the $\Gamma$ point contributed by the $d_{x^{2}-y^{2}}$ orbital. The slight asymmetry between two Ni sites splits the $d_{x^{2}-y^{2}}$ bands along the $\Gamma-M$ path. The further inclusion of interaction in LDA+U formalism will induce an additional electron pocket around the $\Gamma$ point from the $d_{z^{2}}$ anti-bonding state (see SM). In contrast, the bilayer away from the interface, such as the 2nd and 3rd bilayers, shows no charge doping and its band structure closely resembles that of the bulk band structure at ambient pressure, as shown in Fig. 1(c). The $d_{z^{2}}$ bonding state just drops below the Fermi level Zhao and Botana (2025), yielding an electron pocket around the $\Gamma$ point and one hole pocket around the M point. Based on these electronic structures, we conduct model analysis and numerical calculations on correlated states to uncover the scenario of superconductivity within this heterostructure.
[…]
Scenarios of superconductivity in compressed LNO thin films. Although superconductivity has been experimentally achieved in strained LNO ultrathin films, identifying which NiO2 bilayer is responsible remains elusive. Understanding this is crucial for revealing the underlying mechanism of superconductivity, as there is a significant distinction between the electronic structures of NiO2 bilayers near the interface and those further away. To investigate this issue, we construct effective tight-binding models by fitting hopping parameters to ab initio band structures and study electronic instability by varying carrier doping and interaction parameters, based on FRG calculations. As discussed earlier, the Ni $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals dominate the low-energy physics in strained LNO thin films, similar to bulk LNO at both ambient and high pressures Luo et al. (2023); Zhang et al. (2023a); Lechermann et al. (2023a); Sakakibara et al. (2024a); Gu et al. (2025); Yang et al. (2024); Liu et al. (2024). Therefore, a two-orbital tight-binding (TB) model can accurately describe the low-energy band structure Luo et al. (2023); Zhang et al. (2023a); Lechermann et al. (2023a); Sakakibara et al. (2024a); Gu et al. (2025), and the corresponding Hamiltonian is given by,
$\mathcal{H}_{0}=\sum_{ij,\alpha\beta,\sigma}t_{\alpha\beta}^{ij}c_{i\alpha\sigma}^{\dagger}c_{j\beta\sigma}-\mu\sum_{i\alpha\sigma}c_{i\alpha\sigma}^{\dagger}c_{j\alpha\sigma}.$ (1)
Here $i,j=(m,l)$ label the in-plane lattice site ($m$) and top/bottom layer index ($l=t,b$), $\sigma$ labels spin, and $\alpha,\beta=x,z$ represent the Ni orbitals with $x$ denoting the $d_{x^{2}-y^{2}}$ and $z$ the $d_{z^{2}}$ orbital. The chemical potential $\mu$ varies with charge doping. Details of the Hamiltonian matrix are provided in the SM. Compared to the bulk case, the asymmetry between the two Ni sites in the 1st bilayer at the interface breaks the mirror symmetry connecting them, resulting in slightly different onsite energies and in-plane hopping for the two layers. The bilayer slightly away from the interface preserves this mirror symmetry and shares the same TB model with the bulk, though with modified hopping parameters and onsite energies due to strain. For the electronic interactions, we adopt the general multi-orbital Hubbard interactions,
$\mathcal{H}_{\mathrm{I}}$ $=\sum_{i\alpha}Un_{i\alpha\uparrow}n_{i\alpha\downarrow}+\sum_{i,\alpha\neq\beta}J_{P}c_{i\alpha\uparrow}^{\dagger}c_{i\alpha\downarrow}^{\dagger}c_{i\beta\downarrow}c_{i\beta\uparrow}$ (2)
$+\sum_{i,\alpha<\beta,\sigma\sigma^{\prime}}(U^{\prime}n_{i\alpha\sigma}n_{i\beta\sigma^{\prime}}+J_{H}c_{i\alpha\sigma}^{\dagger}c_{i\beta\sigma}c_{i\beta\sigma^{\prime}}^{\dagger}c_{i\alpha\sigma^{\prime}}),$
where $U/U^{\prime}$ is the intra-orbital/inter-orbital Hubbard repulsion, $J_{H}$ is the Hund ’s coupling, and $J_{P}$ is the pair-hopping interaction. The standard Kanamori relations Castellani et al. (1978) $U=U^{\prime}+2J_{H}$ and $J_{H}=J_{P}$ are adopted in this work. To resolve electronic instabilities, we employ the FRG approach, which treats all particle-hole and particle-particle channels on equal footing, providing an unbiased depiction of correlated states from weak to intermediate coupling regimes Wang et al. (2009); Metzner et al. (2012); Platt et al. (2013). As the RG cutoff $\Lambda$ decreases, a divergent flow indicates an instability towards a symmetry-broken state in either the particle-particle or particle-hole channel. Further details of the formalism are provided in the SM.
We begin by examining the electronic instability of the NiO2 bilayer at the interface. Fig. 2(a) presents the orbital-resolved band structure from the effective TB model for this bilayer, and Fig. 2(b) illustrates the corresponding Fermi surfaces. The onsite energy difference for the $d_{z^{2}}$ orbital between the top and bottom layers is approximately 0.23 eV according to Wannier fitting. Compared to the pressurized bulk case, there is an increase in the weight of both the $d_{x^{2}-y^{2}}$ bonding orbital on the $\alpha$ pocket and the $d_{z^{2}}$ antibonding orbital on the $\beta$ pocket. This occurs as the $d_{z^{2}}$ antibonding orbital moves away from the Fermi level, while the antibonding orbital shifts closer to it. In Fig. 2(c), we depict the representative FRG flow for a slightly hole-doped case with a reasonable interaction setting of $U=3$ eV and $J/U=0.1$. The SDW state is dominant from the outset and diverges at a relatively high cutoff, indicating a strong SDW instability, while superconductivity instability is weak. The inset shows the leading eigenvalues in the 2D Brillouin zone (BZ), with divergence occurring at $\bm{q}_{4}\approx(3\pi/4,3\pi/4)$, corresponding to Fermi surface nesting between the $\alpha$ and $\beta$ pockets. To simulate charge doping in realistic systems Zhou et al. (2025), we performed systematic calculations for both hole-doped ($\delta>0$) and electron-doped ($\delta<0$) cases for various Hund’s coupling. The resulting phase diagram, shown in Fig. 2(d), demonstrates a clear dominance of the SDW state with a wavevector on the $\Gamma$-M line across a wide doping regime. In contrast, superconductivity only appears within a narrow range of weak Hund’s coupling with small critical cutoffs. Notably, the leading pairing is $d$-wave for the electron-doped case after the Lifshitz transition (denoted by the black dashed line), where an electron pocket around the $\Gamma$ point arises from the $d_{z^{2}}$ antibonding orbital. Based on these findings, we deduce that the observed superconductivity is unlikely to originate from the interface, consistent with robust superconductivity observed in different samples with different interfaces in experiments Ko et al. (2025); Zhou et al. (2025); Liu et al. (2025).
Figure 3: Band structure, Fermi surfaces and electronic instability for the NiO2 bilayer near the interface. Tight-binding model band structure (a) of the third NiO2 bilayer and corresponding Fermi surface (b) at filling $n=2.7$. (c) Typical FRG flows of leading eigenvalues in SC, SDW and CDW channels with $U=3$ eV, $J=0.1U,n=2.7$ of which leading instability is a $s_{\pm}$-superconductivity with inset figure showing the corresponding superconducting gap on Fermi surfaces. (d) Phase diagram of instability as a function of Hund’s rule coupling strength $J_{H}$ and hole doping level $\delta$ with a fixed $U=3$ eV. The symbols and color bar have the same meaning as in Fig.2. The black dashed line denotes the Lifshitz transition point $\delta_{c}=0.22$ with the appearance of hole pocket around the M point for $\delta>\delta_{c}$.
Then we investigate the instabilities in the NiO2 bilayer slightly away from the interface, which is influenced solely by epitaxial strain. The orbital-resolved band structure is depicted in Fig.3(a), showing two pockets in the absence of charge doping. The disappearance of the $d_{z^{2}}$-bonding $\gamma$ hole pocket is attributed to the reduced onsite energy of $d_{z^{2}}$ orbital due to the compressive strain. However, this hole pocket can emerge through a Lifshitz transition when the hole doping level $\delta$ exceeds $\delta_{c}=0.22$. Fig. 3(b) illustrates the Fermi surfaces in such a case with $\delta=0.3$. Prior to the Lifshitz transition, our FRG calculations show that the system predominantly exhibits SDW instability with either no doping or slight hole doping according to the phase diagram in Fig. 3(d). However, after the Lifshitz transition (black dashed line), superconductivity instability becomes robust and can even persist under strong Hund’s coupling, as shown in Fig. 3(d). A representative FRG flow for the case with a hole doping $\delta=0.3$ is displayed in Fig. 3(c), where $s_{\pm}$-superconductivity is dominant, with the gap function depicted in the inset. This pairing state is primarily attributed to the interlayer pairing, which induces a sign change in superconducting gaps between the bonding and anti-bonding states. This behavior is analogous to that observed in the bulk under pressure Sakakibara et al. (2024a); Gu et al. (2025); Yang et al. (2023a); Liu et al. (2023). These suggest that the emergence of the $\gamma$ hole pocket around the M point has a significant effect on the SDW order and $s_{\pm}$-wave superconductivity, which is related to Fermi surface nesting. Without the $\gamma$ pocket, the nesting between $\alpha$ and $\beta$ pockets contributes to significant SDW fluctuation at $\bm{q}_{4}$, inducing a dominant SDW instability. However, the appearance of the $\gamma$ pocket introduces an additional strong nesting between $\beta$ and $\gamma$ pockets at $\bm{q}_{2}$, competing with SDW fluctuation at $\bm{q}_{4}$. This is apparent from the FRG flow in the SDW channel shown in Fig. 3(c) and this competition suppresses SDW instability. Remarkably, both Fermi surface nestings occur between bonding and anti-bonding pockets and thus generate strong interlayer antiferromagnetic fluctuations (see SM), significantly promoting interlayer pairing and stabilizing $s_{\pm}$-wave superconductivity. Consequently, the observed superconductivity in compressed thin films is likely to occur in NiO2 bilayers slightly away from the interface, where the substrate can contribute to both strain effect and hole doping due to the substitution of La/Pr with Sr Zhou et al. (2025).
[…]
Enhanced SDW instability in stretched LNO thin films. We further investigate the instability of LNO thin films under tensile strain. Experimental results indicate that LNO films on SrTiO3 do not exhibit superconductivity but rather show insulating behavior Bhatt et al. (2025), despite theoretical calculations suggesting that their Fermi surfaces are similar to those of the bulk under pressure Geisler et al. (2025); Zhao and Botana (2025). Fig. 4(a) presents the ab initio band structures in a tetragonal phase with an in-plane lattice constant of $a=3.90$ , consistent with previous calculations Geisler et al. (2025); Zhao and Botana (2025). The corresponding Fermi surfaces from an effective TB model are shown in Fig. 4(b), where all three pockets become more rectangular. It is evident that the $\alpha$ pocket incorporates more $d_{z^{2}}$ orbital, while the contribution of the $d_{x^{2}-y^{2}}$ orbital to the $\beta$ pocket around the M point is also enhanced. These generate enhanced Fermi surface nesting at $\bm{q}_{2}$ and $\bm{q}_{4}$ vectors. More intriguingly, the rectangular $\alpha$ pocket aligns well with the $\gamma$ pocket when shifted by $\bm{q}_{5}=(\pi,\pi)$, indicating strong Fermi surface nesting between them. The bare susceptibility along the high-symmetry path are displayed in Fig. 4(c) for $\delta=0$ and $\delta=\pm 0.1$. A sharp nesting peak around $\bm{q}_{5}$ is dominant, while the peaks at $\bm{q}_{4}$ and $\bm{q}_{2}$ are also prominent. Electron doping can relatively enhance nesting at $\bm{q}_{2,4}$. Due to the competition among multiple types of spin fluctuations, the undoped system does not develop instability for a low $J_{H}/U<0.1$ above the cutoff of $10^{-4}$, as shown in Fig. 4(d). However, when Hund’s coupling is increased, the system exhibits a robust SDW instability with a wavevector of $\bm{q}_{4}$, regardless of hole or electron doping, as shown in Fig. 4(d). The nesting at $\bm{q}_{5}$ involves two bonding pockets and induces interlayer ferromagnetic coupling, which suppresses interlayer spin-singlet pairing. This, together with significant SDW fluctuations at $\bm{q}_{2}$ and $\bm{q}_{4}$, renders superconductivity mostly subdominant. The absence of superconductivity and robust SDW seem to be consistent with the experimental observation of insulating behavior in the resistivity of stretched LNO thin films Bhatt et al. (2025).
[…]
Discussion and conclusion. Our calculations highlight the pivotal role of the hole pocket around the M point in promoting superconductivity in bilayer nickelates under compressive strain, as depicted in the schematic phase diagram in Fig. 5(a). This is intimately related to both its high density of states and bonding state nature. Its appearance introduces an additional Fermi surface nesting with vectors $\bm{q}_{2}$, and leads to frustration of SDW fluctuations, thus destabilizing the SDW instability. Meanwhile, two Fermi surface nestings involving vectors $\bm{q}_{2}$ and $\bm{q}_{4}$ occur between bonding and anti-bonding pockets, generating significant interlayer antiferromagnetic coupling, as shown in Fig. 5(a) and (b). When SDW instabilities are not dominant, their fluctuations will promote interlayer pairing, leading to the emergence of superconductivity. This scenario is consistent with the absence of superconductivity for the NiO2 layer at the interface and under tensile strain. In the former case, electron doping can lead to a Lifshitz transition, where an antibonding electron pocket appears around the $\Gamma$ point. Its nesting with the antibonding $\beta$ pocket contributes to ferromagnetic interlayer coupling, thus suppressing interlayer spin-singlet pairing. In the latter case, the strong nesting at $\bm{q}_{5}$ between two bonding pockets also contributes to ferromagnetic interlayer coupling, hindering interlayer pairing.
The maximum $T_{c}$ observed in strained thin films is approximately 50 K Liu et al. (2025); Zhou et al. (2025), which is significantly lower than that of the bulk under pressure. This value could potentially be increased with future improvements in sample quality. However, our calculations suggest that the $T_{c}$ in strained thin films may consistently remain lower than in the bulk. This is due to the elongation of the out-of-plane Ni-Ni distance under compressive strain and Sr doping, which weakens interlayer antiferromagnetic coupling and interlayer pairing. Therefore, further external pressure can enhance interlayer coupling and raise the transition temperature of strained thin films, similar to the infinite-layer nickelates Wang et al. (2022). Under tensile strain, strong Fermi surface nesting leads to competing SDW fluctuations and the system develops an SDW instability. The Fermi surface nesting can be suppressed by charge doping, potentially allowing superconductivity to emerge when the strong SDW order is diminished.
In summary, we investigate the electronic structures and competing correlated instabilities of strained LNO thin films on substrates to identify the key factors for achieving superconductivity. We find that the compressed NiO2 bilayer at the interface, which is heavily electron-doped from the substrate, is unlikely to exhibit superconductivity. In contrast, the NiO2 bilayer slightly away from the interface displays density-wave instability when undoped or slightly hole-doped. However, when this bilayer is moderately hole-doped, resulting in the emergence of a hole pocket around the M point, two SDW orders compete, destabilizing the SDW instability. Intriguingly, both SDWs exhibit interlayer antiferromagnetic coupling, and their fluctuations promote interlayer pairing, leading to robust $s_{\pm}$-wave superconductivity. This may account for the superconductivity observed in thin films. For the stretched NiO2 bilayer, a robust SDW instability is observed over a wide range of interaction parameters due to significantly enhanced Fermi surface nesting, despite the presence of a hole pocket around the M point. We further discuss the differences between bulk and thin films and strategies to enhance $T_{c}$. Our work establishes a unified scenario for superconductivity in both pressurized bulk and strained thin films of bilayer nickelates.
Acknowledgments. We acknowledge the supports by the Ministry of Science and Technology (Grant No. 2022YFA1403901), National Natural Science Foundation of China (No. 11920101005, No. 11888101, No. 12047503, No. 12322405, No. 12104450) and the New Cornerstone Investigator Program. X.W. is supported by the National Key R&D Program of China (Grant No. 2023YFA1407300) and the National Natural Science Foundation of China (Grants No. 12447103). C.C.L. is supported by the RIKEN TRIP initiative (RIKEN Quantum).
Note added. After completing this work, we became aware of an independent study on superconductivity in thin films Shao et al. (2025). This study focuses on the pairing symmetry in the Amam phase without the $\gamma$ hole pocket around the M point, using an either-band tight-binding model. In contrast, our work systematically explores both superconductivity and charge density wave states for compressed and stretched thin films across various levels of charge doping, highlighting the crucial role of the hole pocket around the M point in promoting superconductivity.